# DPStyler: Dynamic PromptStyler for Source-Free Domain Generalization

### v1

@@ Abstract
Source-Free Domain Generalization (SFDG) aims to develop a model that works for unseen target domains without relying on any source domain. Recent work, PromptStyler, employs text prompts to simulate different distribution shifts in the joint vision-language space, allowing the model to generalize effectively to unseen domains without using any images. However, 1) PromptStyler’s style generation strategy has limitations, as all style patterns are fixed after the first training phase. This leads to the training set in the second training phase being restricted to a limited set of styles. Additionally, 2) the frozen text encoder in PromptStyler result in the encoder’s output varying with the style of the input text prompts, making it difficult for the model to learn domain-invariant features. In this paper, we introduce Dynamic PromptStyler (DPStyler), comprising Style Generation and Style Removal modules to address these issues. The Style Generation module refreshes all styles at every training epoch, while the Style Removal module eliminates variations in the encoder’s output features caused by input styles. Moreover, since the Style Generation module, responsible for generating style word vectors using random sampling or style mixing, makes the model sensitive to input text prompts, we introduce a model ensemble method to mitigate this sensitivity. Extensive experiments demonstrate that our framework outperforms state-of-the-art methods on benchmark datasets.
@@ I Introduction
Deep neural networks achieve optimal performance under the fundamental assumption of independent and identically distributed training and testing data. However, when the target domain exhibits a different distribution from the source domain [1, 2, 3, 4, 5, 6], the effectiveness of these models significantly decreases. To tackle this issue, domain adaptation (DA) methods [7, 8, 9, 10, 11, 12, 13, 14, 15, 16] have been extensively studied. These methods assume access to the target domain and aim to adapt the model to that specific domain. Nonetheless, in many real-world scenarios, the target domain is inaccessible, leading to the domain generalization (DG) paradigm [17, 18, 19, 20, 21, 22, 23, 24, 25]. Domain generalization assumes access to one or more source domains but excludes access to the target domain. The primary goal of domain generalization is to improve the model’s generalization ability to any unseen domain. However, models trained on different source domains often exhibit varying degrees of generalization, making it challenging to determine which domains are beneficial for enhancing the model’s generalization ability. Moreover, the process of gathering and annotating extensive multi-source domain data for training purposes can be financially burdensome and, at times, impractical.
@@ II-B Vision-Language Models
Several works have explored the joint learning of image and text representations [29, 48, 49, 50, 51]. These works utilize image-text pairs for training visual-semantic embeddings, which find application not only in image classification, captioning, and retrieval but also in zero-shot prediction for unseen labels. For instance, CLIP [29] presents a scalable contrastive pre-training approach for learning joint text and image features. Leveraging a vast corpus of $400$ million image-text pairs, CLIP achieves superior zero-shot learning capabilities in classification tasks. Recent studies have demonstrated that leveraging style-specific prompts, such as ‘art,’ ‘sketch,’ or ‘photo’ within their joint vision-language spaces allows for the manipulation of visual features [52, 53, 54, 55]. Specially, Textual Inversion [56] shows that a learnable style word vector for a pseudo-word $\boldsymbol{S}_{*}$ could capture a shared style of images using CLIP with a prompt (eg, “a painting in the style of $\boldsymbol{S}_{*}$”).
@@ III-B Style Removal Module
where $\mathbf{v}$ denotes the input to this module and $\mathbf{W}_{1},\mathbf{W}_{2}$ denote the weights of the two fully connected layers. This module adjusts the feature weights for each channel based on the input feature, which helps to enhance the discriminability of the features. By allocating different levels of attention to each channel, the model can focus more on the domain-invariant aspects of the object itself.
To effectively remove specific style information with this module, we utilize a domain classifier $F_{D}(\cdot)$ and a domain uncertainty loss $\mathcal{L}_{U}$. Firstly, a style prompt template $\mathcal{P}_{i}^{\mathrm{style}}$ (eg, “$\boldsymbol{S}_{i}$-like style”), is initialized. Then we create various style prompts $\{\mathcal{P}_{i}^{\mathrm{style}}\}_{i=1}^{K}$ by replacing $\boldsymbol{S}_{i}$ with the $K$ style word vectors obtained by the style generation module. These style prompts are encoded by text encoder $T(\cdot)$ to get the text features $\{T(\mathcal{P}_{i}^{\mathrm{style}})\in\mathbb{R}^{C}\}_{i=1}^{K}$ corresponding to the $K$ styles. Similar to CLIP’s zero-shot prediction, the input of $F_{D}(\cdot)$ consists of features $R(T(\mathcal{P}_{m}^{\mathrm{content}}\circ\mathcal{P}_{i}^{\mathrm{style}}))\in\mathbb{R}^{C}$ obtained by style remover and the style text features $\{T(\mathcal{P}_{i}^{\mathrm{style}})\in\mathbb{R}^{C}\}_{i=1}^{K}$, the output is probabilities computed by consine similarity. Thus, the probability that a sample has a certain style $\boldsymbol{S}_{j}$ can be written as:
$p_{mij}=\frac{e^{z_{mij}}}{\sum_{k=1}^{K}e^{z_{mik}}}\,.$ (5)
$\mathcal{L}_{U}(\mathcal{P}_{m}^{\mathrm{content}}\circ\mathcal{P}_{i}^{\mathrm{style}})=-\sum_{j=1}^{K}p_{mij}\log p_{mij}\,,$ (6)
where $p_{mij}$ is the probability that sample $\mathcal{P}_{m}^{\mathrm{content}}\circ\mathcal{P}_{i}^{\mathrm{style}}$ has style $\boldsymbol{S}_{i}$ and $K$ is the number of styles created.
III-C Model Ensemble
During our experiments we find that introducing more randomness into the training process makes the model more sensitive to changes in the input text prompts. Different initial text templates will make the model behave differently on each dataset, even if they are semantically similar (eg, “a [class] in a $\boldsymbol{S}_{*}$ style” and “a $\boldsymbol{S}_{*}$ style of a [class]”). This lead us to think about how to reduce the impact of template selection on the model’s effectiveness and harvest a stable and desirable classification result. We propose to integrate the advantages of each template in a model ensemble approach. Specifically, for each text template, we train a model and each model is involved in the inference process, as depicted in the Figure 3. Assuming that $N$ models are trained, each model generates $M$ class scores during inference, yielding a total of $N\times M$ class scores. We select the class corresponding to the largest score value among these class scores as the predicted class for the input image.
III-D Model Training and Inference
@@
IV Experiments
IV-A Evaluation Datasets
IV-B Implementation Details
DPStyler maintains consistent implementation and training with identical configurations across all evaluation datasets. The training process typically requires about $1$ minute when conducted on a single RTX3090 GPU, which is 4.5$\sim$ 4.5$\times$ 4.5 × faster compared to PromptStyler [26]. Further detailed comparisons are available in Section IV-C.
@@ IV-B Implementation Details
DPStyler Ours 93.6 ± 0.2 83.5 ± 0.2 72.5 ± 0.2 48.0 ± 0.1    $\textbf{74.4}^{{\tiny\uparrow 3.6}}$
@@
DPStyler Ours 97.1 ± 0.1 84.0 ± 0.4 82.8 ± 0.1 58.9 ± 0.1    $\textbf{80.7}^{{\tiny\uparrow 3.5}}$
@@
DPStyler Ours 98.4 ± 0.1 83.2 ± 0.1 88.0 ± 0.3 64.7 ± 0.1    $\textbf{83.6}^{{\tiny\uparrow 2.6}}$
TABLE I: Comparison with state-of-the-art domain generalization methods, ZS-CLIP (C) refers to zero-shot CLIP using ‘[class]’ as its text prompt, and ZS-CLIP (PC) indicates zero-shot CLIP using ‘a photo of a [class]’ as its text prompt. Bold indicates the best result, the number next to the ↑ indicates our improvement compared to ZS-CLIP (C), and ‘*’ denotes our reproduced results for PromptStyler.
@@ IV-C Evaluations
Method Venue Accuracy (%) $\uparrow$
Art Painting Cartoon Photo Sketch Avg.
ResNet-50 with pre-trained weights on ImageNet
SelfReg [34] ICCV’2021 87.9 ± 1.0 79.4 ± 1.4 96.8 ± 0.7 78.3 ± 1.2 85.6
GVRT [68] ECCV’2022 87.9 ± 0.3 78.4 ± 1.0 98.2 ± 0.1 75.7 ± 0.4 85.1
ResNet-50 with pre-trained weights from CLIP
ZS-CLIP (C) [29] ICML’2021  88.9 ± 0.0 94.4 ± 0.0 99.3 ± 0.0 79.8 ± 0.0 90.6
ZS-CLIP (PC) [29] ICML’2021  90.8 ± 0.0 93.3 ± 0.0 99.4 ± 0.0 79.3 ± 0.0 90.7
DPStyler Ours 94.0 ± 0.2 95.3 ± 0.0 99.6 ± 0.0 85.5 ± 0.4    $\textbf{93.6}^{{\tiny\uparrow 3.0}}$
ViT-B / 16 with pre-trained weights from CLIP
ZS-CLIP (C) [29] ICML’2021  96.4 ± 0.0 98.7 ± 0.0 99.9 ± 0.0 87.5 ± 0.0 95.6
ZS-CLIP (PC) [29] ICML’2021  97.3 ± 0.0 99.0 ± 0.0 99.9 ± 0.0 88.0 ± 0.0 96.1
DPStyler Ours 97.7 ± 0.0 98.9 ± 0.1 99.9 ± 0.0 91.9 ± 0.5    $\textbf{97.1}^{{\tiny\uparrow 1.5}}$
ViT-L / 14 with pre-trained weights from CLIP
ZS-CLIP (C) [29] ICML’2021  97.2 ± 0.0 99.4 ± 0.0 99.9 ± 0.0 93.9 ± 0.0 97.6
ZS-CLIP (PC) [29] ICML’2021  98.6 ± 0.0 99.5 ± 0.0 99.9 ± 0.0 95.3 ± 0.0 98.3
DPStyler Ours  98.5 ± 0.2 99.6 ± 0.1 100.0 ± 0.0 95.6 ± 0.2    $\textbf{98.4}^{{\tiny\uparrow 0.8}}$
TABLE II: Comparison with SOTA domain generalization methods for per-domain top-1 accuracy on PACS. We repeat each experiment using three different seeds, and report average accuracies with standard errors. ZS-CLIP (C) refers to zero-shot CLIP using ‘[class]’ as its text prompt, and ZS-CLIP (PC) indicates zero-shot CLIP using ‘a photo of a [class]’ as its text prompt. Bold indicates the best result, the number next to the ↑ indicates our improvement compared to ZS-CLIP (C), and ‘*’ denotes our reproduced results for PromptStyler.
Method Venue Accuracy (%) $\uparrow$
Caltech LabelMe SUN09 VOC2007 Avg.
ResNet-50 with pre-trained weights on ImageNet
SelfReg [34] ICCV’2021 96.7 ± 0.4 65.2 ± 1.2 73.1 ± 1.3 76.2 ± 0.7 77.8
GVRT [68] ECCV’2022 98.8 ± 0.1 64.0 ± 0.3 75.2 ± 0.5 77.9 ± 1.0 79.0
ResNet-50 with pre-trained weights from CLIP
ZS-CLIP (C) [29] ICML’2021  99.5 ± 0.0 67.8 ± 0.0 69.5 ± 0.0 80.8 ± 0.0 79.4
ZS-CLIP (PC) [29] ICML’2021  99.8 ± 0.0 69.6 ± 0.0 71.0 ± 0.0 87.7 ± 0.0 82.0
DPStyler Ours 100.0 ± 0.0 73.3 ± 0.9 71.8 ± 0.9 89.0 ± 0.1    $\textbf{83.5}^{{\tiny\uparrow 4.1}}$
ViT-B / 16 with pre-trained weights from CLIP
ZS-CLIP (C) [29] ICML’2021  99.8 ± 0.0 60.9 ± 0.0 69.8 ± 0.0 74.1 ± 0.0 76.2
DPStyler Ours 100.0 ± 0.0 69.1 ± 1.3 76.3 ± 0.2 90.5 ± 0.4    $\textbf{84.0}^{{\tiny\uparrow 7.8}}$
ViT-L / 14 with pre-trained weights from CLIP
ZS-CLIP (C) [29] ICML’2021 100.0 ± 0.0 57.5 ± 0.0 70.5 ± 0.0 82.1 ± 0.0 77.5
DPStyler Ours 100.0 ± 0.0 68.2 ± 0.4 76.4 ± 0.5 88.1 ± 0.3    $\textbf{83.2}^{{\tiny\uparrow 5.7}}$
TABLE III: Comparison with SOTA domain generalization methods for per-domain top-1 accuracy on VLCS. We repeat each experiment using three different seeds, and report average accuracies with standard errors. ZS-CLIP (C) refers to zero-shot CLIP using ‘[class]’ as its text prompt, and ZS-CLIP (PC) indicates zero-shot CLIP using ‘a photo of a [class]’ as its text prompt. Bold indicates the best result, the number next to the ↑ indicates our improvement compared to ZS-CLIP (C), and ‘*’ denotes our reproduced results for PromptStyler.
@@
DPStyler Ours 73.6 ± 0.4 51.2 ± 0.2 82.1 ± 0.6 82.9 ± 0.4    $\textbf{72.5}^{{\tiny\uparrow 5.1}}$
@@
DPStyler Ours 83.6 ± 0.4 67.8 ± 0.5 90.0 ± 0.2 89.7 ± 0.1    $\textbf{82.8}^{{\tiny\uparrow 3.2}}$
@@
DPStyler Ours 88.0 ± 0.4 76.4 ± 0.3 94.3 ± 0.5 93.4 ± 0.3    $\textbf{88.0}^{{\tiny\uparrow 2.3}}$
TABLE IV: Comparison with SOTA domain generalization methods for per-domain top-1 accuracy on OfficeHome. We repeat each experiment using three different seeds, and report average accuracies with standard errors. ZS-CLIP (C) refers to zero-shot CLIP using ‘[class]’ as its text prompt, and ZS-CLIP (PC) indicates zero-shot CLIP using ‘a photo of a [class]’ as its text prompt. Bold indicates the best result, the number next to the ↑ indicates our improvement compared to ZS-CLIP (C), and ‘*’ denotes our reproduced results for PromptStyler.
@@
DPStyler Ours 55.4 ± 0.4 41.5 ± 0.1 56.0 ± 0.1 7.1 ± 0.2 77.5 ± 0.0 50.7 ± 0.1    $\textbf{48.0}^{{\tiny\uparrow 2.1}}$
@@
DPStyler Ours 71.5 ± 0.2 50.6 ± 0.1 66.8 ± 0.3 16.2 ± 0.1 83.7 ± 0.0 64.4 ± 0.1    $\textbf{58.9}^{{\tiny\uparrow 1.5}}$
@@
DPStyler Ours 78.3 ± 0.1 54.6 ± 0.0 73.2 ± 0.1 23.7 ± 0.2 86.7 ± 0.0 71.9 ± 0.2    $\textbf{64.7}^{{\tiny\uparrow 1.6}}$
TABLE V: Comparison with SOTA domain generalization methods for per-domain top-1 accuracy on DomainNet. We repeat each experiment using three different seeds, and report average accuracies with standard errors. ZS-CLIP (C) refers to zero-shot CLIP using ‘[class]’ as its text prompt, and ZS-CLIP (PC) indicates zero-shot CLIP using ‘a photo of a [class]’ as its text prompt. Bold indicates the best result, the number next to the ↑ indicates our improvement compared to ZS-CLIP (C), and ‘*’ denotes our reproduced results for PromptStyler.

### v2

@@ Abstract
Source-Free Domain Generalization (SFDG) aims to develop a model that works for unseen target domains without relying on any source domain. Research in SFDG primarily bulids upon the existing knowledge of large-scale vision-language models and utilizes the pre-trained model’s joint vision-language space to simulate style transfer across domains, thus eliminating the dependency on source domain images. However, how to efficiently simulate rich and diverse styles using text prompts, and how to extract domain-invariant information useful for classification from features that contain both semantic and style information after the encoder, are directions that merit improvement. In this paper, we introduce Dynamic PromptStyler (DPStyler), comprising Style Generation and Style Removal modules to address these issues. The Style Generation module refreshes all styles at every training epoch, while the Style Removal module eliminates variations in the encoder’s output features caused by input styles. Moreover, since the Style Generation module, responsible for generating style word vectors using random sampling or style mixing, makes the model sensitive to input text prompts, we introduce a model ensemble method to mitigate this sensitivity. Extensive experiments demonstrate that our framework outperforms state-of-the-art methods on benchmark datasets. Code is available here.
@@ I Introduction
Deep neural networks achieve optimal performance under the fundamental assumption of independent and identically distributed training and testing data. However, when the target domain exhibits a different distribution from the source domain [1, 2, 3, 4, 5, 6], the effectiveness of these models significantly decreases. To tackle this issue, domain adaptation (DA) methods [7, 8, 9, 10, 11, 12, 13, 14, 15, 16] have been extensively studied. These methods assume access to the target domain and aim to adapt the model to that specific domain. Nonetheless, in many real-world scenarios, the target domain is inaccessible, leading to the domain generalization (DG) paradigm [17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. Domain generalization assumes access to one or more source domains but excludes access to the target domain. The primary goal of domain generalization is to improve the model’s generalization ability to any unseen domain. However, models trained on different source domains often exhibit varying degrees of generalization, making it challenging to determine which domains are beneficial for enhancing the model’s generalization ability. Moreover, the process of gathering and annotating extensive multi-source domain data for training purposes can be financially burdensome and, at times, impractical.
@@ II-B Vision-Language Models
Several works have explored the joint learning of image and text representations [30, 49, 50, 51, 52, 53]. These works utilize image-text pairs for training visual-semantic embeddings, which find application not only in image classification, captioning, and retrieval but also in zero-shot prediction for unseen labels. For instance, CLIP [30] presents a scalable contrastive pre-training approach for learning joint text and image features. Leveraging a vast corpus of $400$ million image-text pairs, CLIP achieves superior zero-shot learning capabilities in classification tasks. Recent studies have demonstrated that leveraging style-specific prompts, such as ‘art,’ ‘sketch,’ or ‘photo’ within their joint vision-language spaces allows for the manipulation of visual features [54, 55, 56, 57]. Specially, Textual Inversion [58] shows that a learnable style word vector for a pseudo-word $\boldsymbol{S}_{*}$ could capture a shared style of images using CLIP with a prompt (eg, “a painting in the style of $\boldsymbol{S}_{*}$”).
@@ III-B Style Removal Module
where $\mathbf{v}$ denotes the input to this module and $\mathbf{W}_{1},\mathbf{W}_{2}$ denote the weights of the two fully connected layers. The features output by the image encoder contain both semantic and style information, and our goal is to decouple these two types of information, retaining the semantic information that is useful for classification. So the objective of Style-SE Net is to highlight semantic information and reduce the influence of style information by assigning different weights to different channels of the features, which means channels that represent style information will be given smaller weights.
@@
To effectively remove specific style information with this module, we utilize a domain classifier $F_{D}(\cdot)$ and a domain uncertainty loss $\mathcal{L}_{U}$ to guide the training of Style-SE Net. Firstly, a style prompt template $\mathcal{P}_{i}^{\mathrm{style}}$ (eg, “$\boldsymbol{S}_{i}$-like style”), is initialized. Then we create various style prompts $\{\mathcal{P}_{i}^{\mathrm{style}}\}_{i=1}^{K}$ by replacing $\boldsymbol{S}_{i}$ with the $K$ style word vectors obtained by the style generation module. These style prompts are encoded by text encoder $T(\cdot)$ to get the text features $\{T(\mathcal{P}_{i}^{\mathrm{style}})\in\mathbb{R}^{C}\}_{i=1}^{K}$ corresponding to the $K$ styles. Similar to CLIP’s zero-shot prediction, the input of $F_{D}(\cdot)$ consists of features $R(T(\mathcal{P}_{m}^{\mathrm{content}}\circ\mathcal{P}_{i}^{\mathrm{style}}))\in\mathbb{R}^{C}$ obtained by style remover and the style text features $\{T(\mathcal{P}_{i}^{\mathrm{style}})\in\mathbb{R}^{C}\}_{i=1}^{K}$, the output is probabilities computed by consine similarity. Thus, the probability that a sample has a certain style $\boldsymbol{S}_{j}$ can be written as:
$p_{mij}=\frac{e^{z_{mij}}}{\sum_{k=1}^{K}e^{z_{mik}}}\,.$ (5)
$\mathcal{L}_{U}(\mathcal{P}_{m}^{\mathrm{content}}\circ\mathcal{P}_{i}^{\mathrm{style}})=\sum_{j=1}^{K}p_{mij}\log p_{mij}\,,$ (6)
where $p_{mij}$ is the probability that sample $\mathcal{P}_{m}^{\mathrm{content}}\circ\mathcal{P}_{i}^{\mathrm{style}}$ has style $\boldsymbol{S}_{i}$ and $K$ is the number of styles created. The domain uncertainty loss aims to diminish the specific expression of style information in the features, thereby suppressing the channels associated with style information and implicitly highlighting the domain-invariant information.
TABLE I: Comparison with state-of-the-art domain generalization methods, ZS-CLIP (C) refers to zero-shot CLIP using ‘[class]’ as its text prompt, and ZS-CLIP (PC) indicates zero-shot CLIP using ‘a photo of a [class]’ as its text prompt. Bold indicates the best result. ‘*’ denotes our reproduced results for PromptStyler, and ‘†’ denotes that the paper corresponding to the method does not give the standard errors.
@@
Text Regularization [63] PRCV’2023  87.2 ± 0.0 80.3 ± 0.0 70.4 ± 0.4 44.0 ± 0.0 70.5
Model Ratatouille [64] ICML’2023 89.8 ± 0.0 78.3 ± 0.0 73.5 ± 0.0 47.7 ± 0.0 72.3
@@
DPStyler Ours 93.6 ± 0.2 83.5 ± 0.2 72.5 ± 0.2 48.0 ± 0.1 74.4
@@
Text Regularization [63] PRCV’2023  95.9 ± 0.0 83.0 ± 0.0 82.3 ± 0.0 57.9 ± 0.0 79.8
DPStyler Ours 97.1 ± 0.1 84.0 ± 0.4 82.8 ± 0.1 58.9 ± 0.1 80.7
@@
DPStyler Ours 98.4 ± 0.1 83.2 ± 0.1 88.0 ± 0.3 64.7 ± 0.1 83.6
III-C Model Ensemble
During our experiments we find that introducing more randomness into the training process makes the model more sensitive to changes in the input text prompts. Different initial text templates will make the model behave differently on each dataset, even if they are semantically similar (eg, “a [class] in a $\boldsymbol{S}_{*}$ style” and “a $\boldsymbol{S}_{*}$ style of a [class]”). This lead us to think about how to reduce the impact of template selection on the model’s effectiveness and harvest a stable and desirable classification result. We propose to integrate the advantages of each template in a model ensemble approach [66]. Specifically, for each text template, we train a model and each model is involved in the inference process, as depicted in the Figure 3. Assuming that $N$ models are trained, each model generates $M$ class scores during inference, yielding a total of $N\times M$ class scores. We select the class corresponding to the largest score value among these class scores as the predicted class for the input image.
III-D Model Training and Inference
TABLE II: Comparison with SOTA domain generalization methods for per-domain top-1 accuracy on PACS. We repeat each experiment using three different seeds, and report average accuracies with standard errors. ZS-CLIP (C) refers to zero-shot CLIP using ‘[class]’ as its text prompt, and ZS-CLIP (PC) indicates zero-shot CLIP using ‘a photo of a [class]’ as its text prompt. Bold indicates the best result. ‘*’ denotes our reproduced results for PromptStyler, and ‘†’ denotes that the paper corresponding to the method does not give the standard errors.
Method Venue Accuracy (%) $\uparrow$
Art Painting Cartoon Photo Sketch Avg.
ResNet-50 with pre-trained weights on ImageNet
GVRT [62] ECCV’2022  87.9 ± 0.3 78.4 ± 1.0 98.2 ± 0.1 75.7 ± 0.4 85.1
SAGM [68] CVPR’2023  85.3 ± 0.0 80.6 ± 0.0 95.9 ± 0.0 80.1 ± 0.0 85.5
SelfReg [35] ICCV’2021 87.9 ± 1.0 79.4 ± 1.4 96.8 ± 0.7 78.3 ± 1.2 85.6
M2-CL [69] TAI’2024  87.3 ± 0.0 81.8 ± 0.0 98.5 ± 0.0 76.3 ± 0.0 86.0
JPDA [70] PR’2023 90.6 ± 0.0 83.6 ± 0.0 99.0 ± 0.0 82.8 ± 0.0 89.0
Model Ratatouille [64] ICML’2023 90.6 ± 0.0 84.7 ± 0.0 98.8 ± 0.0 85.0 ± 0.0 89.8
ResNet-50 with pre-trained weights from CLIP
ZS-CLIP (C) [30] ICML’2021  88.9 ± 0.0 94.4 ± 0.0 99.3 ± 0.0 79.8 ± 0.0 90.6
ZS-CLIP (PC) [30] ICML’2021  90.8 ± 0.0 93.3 ± 0.0 99.4 ± 0.0 79.3 ± 0.0 90.7
DPStyler Ours 94.0 ± 0.2 95.3 ± 0.0 99.6 ± 0.0 85.5 ± 0.4 93.6
ViT-B / 16 with pre-trained weights from CLIP
ZS-CLIP (C) [30] ICML’2021  96.4 ± 0.0 98.7 ± 0.0 99.9 ± 0.0 87.5 ± 0.0 95.6
ZS-CLIP (PC) [30] ICML’2021  97.3 ± 0.0 99.0 ± 0.0 99.9 ± 0.0 88.0 ± 0.0 96.1
DPStyler Ours 97.7 ± 0.0 98.9 ± 0.1 99.9 ± 0.0 91.9 ± 0.5 97.1
ViT-L / 14 with pre-trained weights from CLIP
ZS-CLIP (C) [30] ICML’2021  97.2 ± 0.0 99.4 ± 0.0 99.9 ± 0.0 93.9 ± 0.0 97.6
ZS-CLIP (PC) [30] ICML’2021  98.6 ± 0.0 99.5 ± 0.0 99.9 ± 0.0 95.3 ± 0.0 98.3
DPStyler Ours  98.5 ± 0.2 99.6 ± 0.1 100.0 ± 0.0 95.6 ± 0.2 98.4
IV Experiments
IV-A Evaluation Datasets
IV-B Implementation Details
DPStyler maintains consistent implementation and training with identical configurations across all evaluation datasets. The training process typically requires about $1$ minute when conducted on a single RTX3090 GPU, which is 4.5$\sim$ 4.5$\times$ 4.5 × faster compared to PromptStyler [27]. Further detailed comparisons are available in Section III.
@@ IV-C Evaluations
TABLE III: Comparison with SOTA domain generalization methods for per-domain top-1 accuracy on VLCS. We repeat each experiment using three different seeds, and report average accuracies with standard errors. ZS-CLIP (C) refers to zero-shot CLIP using ‘[class]’ as its text prompt, and ZS-CLIP (PC) indicates zero-shot CLIP using ‘a photo of a [class]’ as its text prompt. Bold indicates the best result. ‘*’ denotes our reproduced results for PromptStyler, and ‘†’ denotes that the paper corresponding to the method does not give the standard errors.
Method Venue Accuracy (%) $\uparrow$
Caltech LabelMe SUN09 VOC2007 Avg.
ResNet-50 with pre-trained weights on ImageNet
SAGM [68] CVPR’2023  97.9 ± 0.0 65.0 ± 0.0 70.6 ± 0.0 76.3 ± 0.0 77.4
SelfReg [35] ICCV’2021 96.7 ± 0.4 65.2 ± 1.2 73.1 ± 1.3 76.2 ± 0.7 77.8
M2-CL [69] TAI’2024  98.2 ± 0.0 63.9 ± 0.0 72.4 ± 0.0 76.7 ± 0.0 77.8
Model Ratatouille [64] ICML’2023 99.3 ± 0.0 60.4 ± 0.0 73.9 ± 0.0 79.5 ± 0.0 78.3
GVRT [62] ECCV’2022  98.8 ± 0.1 64.0 ± 0.3 75.2 ± 0.5 77.9 ± 1.0 79.0
ResNet-50 with pre-trained weights from CLIP
ZS-CLIP (C) [30] ICML’2021  99.5 ± 0.0 67.8 ± 0.0 69.5 ± 0.0 80.8 ± 0.0 79.4
ZS-CLIP (PC) [30] ICML’2021  99.8 ± 0.0 69.6 ± 0.0 71.0 ± 0.0 87.7 ± 0.0 82.0
DPStyler Ours 100.0 ± 0.0 73.3 ± 0.9 71.8 ± 0.9 89.0 ± 0.1 83.5
ViT-B / 16 with pre-trained weights from CLIP
ZS-CLIP (C) [30] ICML’2021  99.8 ± 0.0 60.9 ± 0.0 69.8 ± 0.0 74.1 ± 0.0 76.2
DPStyler Ours 100.0 ± 0.0 69.1 ± 1.3 76.3 ± 0.2 90.5 ± 0.4 84.0
ViT-L / 14 with pre-trained weights from CLIP
ZS-CLIP (C) [30] ICML’2021 100.0 ± 0.0 57.5 ± 0.0 70.5 ± 0.0 82.1 ± 0.0 77.5
DPStyler Ours 100.0 ± 0.0 68.2 ± 0.4 76.4 ± 0.5 88.1 ± 0.3 83.2
@@
TABLE IV: Comparison with SOTA domain generalization methods for per-domain top-1 accuracy on OfficeHome. We repeat each experiment using three different seeds, and report average accuracies with standard errors. ZS-CLIP (C) refers to zero-shot CLIP using ‘[class]’ as its text prompt, and ZS-CLIP (PC) indicates zero-shot CLIP using ‘a photo of a [class]’ as its text prompt. Bold indicates the best result. ‘*’ denotes our reproduced results for PromptStyler, and ‘†’ denotes that the paper corresponding to the method does not give the standard errors.
@@
DPStyler Ours 73.6 ± 0.4 51.2 ± 0.2 82.1 ± 0.6 82.9 ± 0.4 72.5
@@
DPStyler Ours 83.6 ± 0.4 67.8 ± 0.5 90.0 ± 0.2 89.7 ± 0.1 82.8
@@
DPStyler Ours 88.0 ± 0.4 76.4 ± 0.3 94.3 ± 0.5 93.4 ± 0.3 88.0
TABLE V: Comparison with SOTA domain generalization methods for per-domain top-1 accuracy on DomainNet. We repeat each experiment using three different seeds, and report average accuracies with standard errors. ZS-CLIP (C) refers to zero-shot CLIP using ‘[class]’ as its text prompt, and ZS-CLIP (PC) indicates zero-shot CLIP using ‘a photo of a [class]’ as its text prompt. Bold indicates the best result. ‘*’ denotes our reproduced results for PromptStyler, and ‘†’ denotes that the paper corresponding to the method does not give the standard errors.
@@
DPStyler Ours 55.4 ± 0.4 41.5 ± 0.1 56.0 ± 0.1 7.1 ± 0.2 77.5 ± 0.0 50.7 ± 0.1 48.0
@@
DPStyler Ours 71.5 ± 0.2 50.6 ± 0.1 66.8 ± 0.3 16.2 ± 0.1 83.7 ± 0.0 64.4 ± 0.1 58.9
@@
DPStyler Ours 78.3 ± 0.1 54.6 ± 0.0 73.2 ± 0.1 23.7 ± 0.2 86.7 ± 0.0 71.9 ± 0.2 64.7